The length of a rectangle is 10 inches more than the width. the perimeter of the rectangle is 112 inches. Can you please help me find what the width of the rectangle is?
Answer by mananth(16946)

let width be x
length = x+10
..
Perimeter = 2*(L+W)
2(x+x+10)=112
2(2x+10)=112
4x+20 = 112
4x=92
/4
x= 92/4
x 23 inches the width
length = 23+10 = 33 inches.
...
CHECK
33+33+23+23=112
...
m.ananth@hotmail.ca